# Riemannian Flow Matching Policy for Robot Motion Learning

### v1

@@ Abstract
We introduce Riemannian Flow Matching Policies (RFMP), a novel model for learning and synthesizing robot visuomotor policies. RFMP leverages the efficient training and inference capabilities of flow matching methods. By design, RFMP inherits the strengths of flow matching: the ability to encode high-dimensional multimodal distributions, commonly encountered in robotic tasks, and a very simple and fast inference process. We demonstrate the applicability of RFMP to both state-based and vision-conditioned robot motion policies. Notably, as the robot state resides on a Riemannian manifold, RFMP inherently incorporates geometric awareness, which is crucial for realistic robotic tasks. To evaluate RFMP, we conduct two proof-of-concept experiments, comparing its performance against Diffusion Policies. Although both approaches successfully learn the considered tasks, our results show that RFMP provides smoother action trajectories with significantly lower inference times.
@@ II-A Riemannian manifolds
Let us imagine a flexible and curved surface like a globe. A smooth manifold, mathematically denoted by $\mathcal{M}$, can be intuitively conceptualized as a smaller patch on that surface. Locally, this patch looks flat, therefore resembling the Euclidean space $\mathbb{R}^{d}$ [20, 21]. But unlike the entire globe, the whole manifold may be curved or twisted, preventing it from being entirely flat. The smoothness of the manifold allows us to define directions and rates of change at each point, leading to tangent vectors in $\mathbb{R}^{d}$. The set of tangent vectors of all curves at $\bm{x}\in\mathcal{M}$ forms a $d$-dimensional affine subspace of $\mathbb{R}^{d}$, known as the tangent space $\mathcal{T}_{\bm{x}}\mathcal{M}$ of $\mathcal{M}$ at $\bm{x}$. The collection of all such tangent spaces is called the tangent bundle $\mathcal{T}\mathcal{M}=\bigcup_{\bm{x}\in\mathcal{M}}\left\{(\bm{x},\bm{u})|\bm{u}\in\mathcal{T}_{x}\mathcal{M}\right\}$ It is possible to endow a smooth manifold $\mathcal{M}$ with a Riemannian metric, which is a family of inner products $g_{\bm{x}}:\mathcal{T}_{\bm{x}}\mathcal{M}\times\mathcal{T}_{\bm{x}}\mathcal{M}\rightarrow\mathbb{R}$ associated to each point $\bm{x}\in\mathcal{M}$. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold endowed with a Riemannian metric $g$, that is a family of inner products $g_{\bm{x}}:\mathcal{T}_{\bm{x}}\mathcal{M}\times\mathcal{T}_{\bm{x}}\mathcal{M}\rightarrow\mathbb{R}$ associated to each point $\bm{x}\in\mathcal{M}$ [21].
To operate with Riemannian manifolds, we can leverage their Euclidean tangent spaces and resort to mappings back and forth between $\mathcal{T}_{\bm{x}}\mathcal{M}$ and $\mathcal{M}$, using the exponential and logarithmic maps. Specifically, the exponential map $\text{Exp}_{\bm{x}}\left(\bm{u}\right):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{M}$ maps a point $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to a point $\bm{y}$ on the manifold, so that it lies on the geodesic starting at $\bm{x}$ in the direction $\bm{u}$ and such that the geodesic distance $d_{\mathcal{M}}(\bm{x},\bm{y})=d_{\mathbb{R}}(\bm{x},\bm{u})$. The inverse operation is the logarithmic map $\text{Log}_{\bm{x}}\left(\bm{y}\right):\mathcal{M}\to\mathcal{T}_{\bm{x}}\mathcal{M}$ Finally, the parallel transport $\Gamma_{\bm{x}\rightarrow\bm{y}}(\bm{u}):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{T}_{\bm{y}}\mathcal{M}$ describes how elements of $\mathcal{M}$ can be transported along curves on $\mathcal{M}$ while maintaining their intrinsic geometric properties. This allows us to operate elements lying on different tangent spaces.
@@ II-B Flow Matching
where $\bm{\theta}$ are the learnable parameters, $t\sim\mathcal{U}[0,1]$, and $\bm{x}\sim p_{t}(\bm{x})$. Unfortunately, the objective in (2) is intractable since we actually do not have prior knowledge about $p_{t}$ and $u_{t}$. Lipman et al. [8] showed that by defining a conditional probability path $p_{t}(\bm{x}|\bm{z})$ (and consequently a conditional vector field $u_{t}(\bm{x}|\bm{z})$, it is possible to obtain a tractable conditional flow matching (CFM) loss,
@@
We now can design the vector field $u_{t}(\bm{x}|\bm{z})$ by leveraging the ODE associated with the conditional flow $d\phi_{t}(\bm{x})/dt=\dot{\bm{x}}_{t}=u_{t}(\bm{x}|\bm{z})$. This means that computing the vector field $u_{t}(\bm{x}|\bm{z})$ corresponds to compute the time derivative of (7). Finally, the choice of the base distribution $p_{0}$ generally depends on the problem at hand. One could directly define the base density as a uniform distribution over $\mathcal{M}$ as in [19, 11], but it is also possible to use Riemannian or wrapped Gaussian distributions.
@@ III-A RFMP training
2       Sample time step $t\sim\mathcal{U}$. Sample training example $\bm{a}_{1}\sim p$, and noise $\bm{a}_{0}\sim q$. Sample observation vector $\bm{o}$. Compute target vector field $\dot{\bm{a}}_{t}=u_{t}(\bm{a}|\bm{a}_{1})$ based on the geodesic flow (7). Evaluate $\mathcal{L}_{\text{RFMP}}(\bm{\theta})$ (8). Update parameters $\bm{\theta}$.

### v2

@@ Abstract
We introduce Riemannian Flow Matching Policies (RFMP), a novel model for learning and synthesizing robot sensorimotor policies. RFMP leverages the efficient training and inference capabilities of flow matching methods. By design, RFMP inherits the strengths of flow matching: the ability to encode high-dimensional multimodal distributions, commonly encountered in robotic tasks, and a very simple and fast inference process. We demonstrate the applicability of RFMP to both state-based and vision-conditioned robot motion policies. Notably, as the robot state resides on a Riemannian manifold, RFMP inherently incorporates geometric awareness, which is crucial for realistic robotic tasks. To evaluate RFMP, we conduct two proof-of-concept experiments, comparing its performance against Diffusion Policies. Although both approaches successfully learn the considered tasks, our results show that RFMP provides smoother action trajectories with significantly lower inference times.
@@ II-A Riemannian manifolds
Let us imagine a flexible and smoothly-curved surface like a globe. A smooth manifold, mathematically denoted by $\mathcal{M}$, can be intuitively conceptualized as the globe surface. Locally, a small patch of this manifold looks flat, therefore resembling the Euclidean space $\mathbb{R}^{d}$ [20, 21]. The smoothness of the manifold allows us to define directions and rates of change at each point, leading to tangent vectors in $\mathbb{R}^{d}$. The set of tangent vectors of all curves at $\bm{x}\in\mathcal{M}$ forms a vector space isomorphic to $\mathbb{R}^{d}$, known as the tangent space $\mathcal{T}_{\bm{x}}\mathcal{M}$ of $\mathcal{M}$ at $\bm{x}$. The collection of all such tangent spaces is called the tangent bundle $\mathcal{T}\mathcal{M}=\bigcup_{\bm{x}\in\mathcal{M}}\left\{(\bm{x},\bm{u})|\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}\right\}$ It is possible to endow a smooth manifold $\mathcal{M}$ with a Riemannian metric, which is a family of inner products $g_{\bm{x}}:\mathcal{T}_{\bm{x}}\mathcal{M}\times\mathcal{T}_{\bm{x}}\mathcal{M}\rightarrow\mathbb{R}$ associated to each point $\bm{x}\in\mathcal{M}$. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold endowed with a Riemannian metric $g$, that is a family of inner products $g_{\bm{x}}:\mathcal{T}_{\bm{x}}\mathcal{M}\times\mathcal{T}_{\bm{x}}\mathcal{M}\rightarrow\mathbb{R}$ associated to each point $\bm{x}\in\mathcal{M}$ [21].
To operate with Riemannian manifolds, we can leverage their Euclidean tangent spaces and resort to mappings back and forth between $\mathcal{T}_{\bm{x}}\mathcal{M}$ and $\mathcal{M}$, using the exponential and logarithmic maps. Specifically, the exponential map $\text{Exp}_{\bm{x}}\left(\bm{u}\right):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{M}$ maps a point $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to a point $\bm{y}$ on the manifold, so that it lies on the geodesic starting at $\bm{x}$ in the direction $\bm{u}$ and such that the geodesic distance $d_{\mathcal{M}}(\bm{x},\bm{y})=d_{\mathbb{R},g_{\bm{x}}}(\bm{x},\bm{u})$. The inverse operation is the logarithmic map $\text{Log}_{\bm{x}}\left(\bm{y}\right):\mathcal{M}\to\mathcal{T}_{\bm{x}}\mathcal{M}$ Finally, the parallel transport $\Gamma_{\bm{x}\rightarrow\bm{y}}(\bm{u}):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{T}_{\bm{y}}\mathcal{M}$ describes how elements of $\mathcal{M}$ can be transported along curves on $\mathcal{M}$ while maintaining their intrinsic geometric properties. This allows us to operate elements lying on different tangent spaces.
@@ II-B Flow Matching
where $\bm{\theta}$ are the learnable parameters, $t\sim\mathcal{U}[0,1]$, and $\bm{x}\sim p_{t}(\bm{x})$. Unfortunately, the objective in (2) is intractable since we actually do not have prior knowledge about $p_{t}$ and $u_{t}$. Lipman et al. [8] showed that by defining a conditional probability path $p_{t}(\bm{x}|\bm{z})$ (and consequently a conditional vector field $u_{t}(\bm{x}|\bm{z})$), it is possible to obtain a tractable conditional flow matching (CFM) loss,
@@
We now can design the vector field $u_{t}(\bm{x}|\bm{z})$ by leveraging the ODE associated with the conditional flow $d\phi_{t}(\bm{x})/dt~{}=~{}\dot{\bm{x}}_{t}~{}=~{}u_{t}(\bm{x}|\bm{z})$. This means that computing the vector field $u_{t}(\bm{x}|\bm{z})$ corresponds to compute the time derivative of (7). Finally, the choice of the base distribution $p_{0}$ generally depends on the problem at hand. One could directly define the base density as a uniform distribution over $\mathcal{M}$ as in [19, 11], but it is also possible to use Riemannian or wrapped Gaussian distributions, as we do next.
@@ III-A RFMP training
2       Sample time step $t\sim\mathcal{U}$.
3       Sample training example $\bm{a}_{1}\sim p$, and noise $\bm{a}_{0}\sim q$.
4       Sample observation vector $\bm{o}$.
5       Compute target vector field $\dot{\bm{a}}_{t}=u_{t}(\bm{a}|\bm{a}_{1})$ based on the geodesic flow (7).
6       Evaluate $\mathcal{L}_{\text{RFMP}}(\bm{\theta})$ (8).
7       Update parameters $\bm{\theta}$.